# BlindGuard: Safeguarding LLM-based Multi-Agent Systems under Unknown Attacks

### v1

@@ Abstract
The security of LLM-based multi-agent systems (MAS) is critically threatened by propagation vulnerability, where malicious agents can distort collective decision-making through inter-agent message interactions. While existing supervised defense methods demonstrate promising performance, they may be impractical in real-world scenarios due to their heavy reliance on labeled malicious agents to train a supervised malicious detection model. To enable practical and generalizable MAS defenses, in this paper, we propose BlindGuard, an unsupervised defense method that learns without requiring any attack-specific labels or prior knowledge of malicious behaviors. To this end, we establish a hierarchical agent encoder to capture individual, neighborhood, and global interaction patterns of each agent, providing a comprehensive understanding for malicious agent detection. Meanwhile, we design a corruption-guided detector that consists of directional noise injection and contrastive learning, allowing effective detection model training solely on normal agent behaviors. Extensive experiments show that BlindGuard effectively detects diverse attack types (i.e., prompt injection, memory poisoning, and tool attack) across MAS with various communication patterns while maintaining superior generalizability compared to supervised baselines. The code is available at: https://github.com/MR9812/BlindGuard.
Introduction
@@
Rapid advancements in large language models (LLMs) have significantly improved their performance in various domains, including task planning (Kannan, Venkatesh, and Min 2024), mathematical reasoning (Lei et al. 2024), and scientific simulations (Zheng et al. 2023). By incorporating modular extensions such as memory (Zhang et al. 2024), tool usage (Masterman et al. 2024), and role-playing capabilities (Kim et al. 2024), LLM-based autonomous agents have expanded their applicability, enabling more dynamic and interactive functionalities (Li et al. 2024b). Building upon these advances, multi-agent systems (MAS) further amplify these benefits by facilitating collaborative interactions among specialized agents (Guo et al. 2024). Recent studies have shown that MAS outperform individual agents in more complex tasks such as problem solving (Li et al. 2025b), embodied actions (Zhao et al. 2024), and social simulations (Zhao et al. 2023). However, the increased reliance on inter-agent communication introduces additional risks in security and controllability, necessitating robust frameworks to safeguard sensitive data and regulate information flow.
@@
Figure 1: Comparison of supervised vs. unsupervised graph anomaly detection-based defense paradigms in MAS.
@@
Security studies (Andriushchenko et al. 2025; Gan et al. 2024; He et al. 2025) have identified significant vulnerabilities in external components of LLM-based agents, including tool interfaces (Zhan et al. 2024a) and memory modules (Chen et al. 2024). Beyond these risks at the single-agent level, the transition to MAS brings additional vulnerabilities caused by inter-agent interactions (Yu et al. 2025). Specifically, the misleading message generated by a few malicious agents can propagate through collaborative reasoning, negatively affecting how agents make collective decisions. Such propagation vulnerability makes MAS susceptible to attacks such as prompt injection through compromised agents, misinformation propagation, and emergent malicious coordination (Yu et al. 2024; Wang et al. 2025).
Aiming to mitigate the propagation vulnerability, graph-based defense provides a promising solution against adversarial attacks in MAS (Zhang et al. 2025b). As a semantically structured data format, graphs can naturally model both the functional roles of individual agents and their inter-agent interactions in MAS (Shen et al. 2025; Liu et al. 2025; Li et al. 2025a). Building upon the graph-based MAS formulation, G-Safeguard (Wang et al. 2025) integrates a detection-remediation framework to effectively safeguard MAS. As demonstrated in Figure 1a, G-Safeguard employs a supervised graph anomaly detection (GAD) model as its core component to identify malicious agents, afterwards applies an edge pruning-based remediation strategy to isolate and suppress the influence of compromised agents.
@@
Despite its impressive defensive performance, G-Safeguard, or other supervised GAD-based approaches, may be impractical in real-world scenarios due to their heavy reliance on labeled malicious agents. Specifically, the supervised paradigm requires labeled instances of actual malicious agents associated with a particular attack type to train a type-specific binary GAD model, which limits its availability and generalizability. On the one hand, adversarial attack behaviors in real-world scenarios are sparse and often purposefully camouflaged, making it difficult to obtain well-annotated malicious agents for supervised training. This inaccessibility of labeled data significantly undermines the availability of supervised GAD-based methods in real-world MAS deployments. On the other hand, real-world MAS face diverse and evolving adversarial attacks, while conventional binary GAD models are typically trained to detect a specific type of malicious behavior (Wang et al. 2025). Such a single-purpose design limits their generalizability and makes them ineffective for detecting novel or unseen attack patterns in complex environments. These limitations raise a critical research question: Can we design a defense framework for MAS without relying on labeled attack agents?
@@
To answer the above question, unsupervised GAD offers a promising solution, where GAD models learn to identify irregular patterns without the supervision of labeled anomalous instances (Ding et al. 2019; Ma et al. 2022; Qiao and Pang 2023). As illustrated in Figure 1b, based on unsupervised GAD techniques, we can train a detector to identify malicious agents associated with multiple attack types using only normal MAS interaction data, thereby alleviating the limitations in availability and generalizability. Nevertheless, directly applying existing unsupervised GAD methods, which are not specifically designed for MAS scenarios, may lead to suboptimal performance due to the following gaps. Gap 1 - Limited multi-level contextual awareness: Identifying malicious agents in MAS requires integrating information across multiple levels, including individual behaviors, local neighborhoods, and global system dynamics. However, most existing GAD methods (Qiao and Pang 2023; Liu et al. 2021; Pan et al. 2023) primarily focus on local properties (e.g., local affinity and ego-neighbor similarity), lacking the system-level understanding needed for multi-agent interaction networks. Gap 2 - Misalignment of Anomalous Behavior Assumptions: Most unsupervised GAD methods assume anomalies manifest through structural deviations (e.g., low homophily (Qiao and Pang 2023) or rare connectivity patterns (Liu et al. 2021)). In contrast, malicious agents in MAS often exhibit semantic anomalies (such as deceptive intent or information poisoning (Yu et al. 2024)) that do not well match these assumptions.
@@
To fill the gaps, in this paper, we propose a novel defense method for MAS, termed BlindGuard, that can be trained without any labeled malicious data or prior knowledge of attack strategies. In BlindGuard, we design a MAS-specific unsupervised GAD model for malicious agent identification, followed by an edge pruning-based remediation module to suppress adversarial propagation. To bridge Gap 1, we introduce a hierarchical agent encoder to incorporate the information of individual agent features, local neighborhood aggregation, and global system context simultaneously. As a result, the encoder captures comprehensive representations of agents to support accurate malicious agent detection. To mitigate Gap 2, we propose a corruption-guided attack detector for agent abnormality estimation. To train the detector, we simulate the malicious behaviors via semantic-level corruption, which is utilized to optimize the detector via a supervised contrastive learning objective. The training of BlindGuard only requires a small amount of normal MAS interaction data, and the learned model can generalize to various types of attacks.
To sum up, the contributions of this paper are three-fold:
@@
  • Scenario. We investigate the scenario of MAS safeguarding without relying on labeled attack data or prior knowledge of attacks, which is more practical and applicable to real-world MAS with limited supervision.
@@
  • Experiments. We extensively evaluate BlindGuard under rigorous real-world conditions with different types of attack. Through comprehensive testing on 4 MAS interaction patterns with 3 attack strategies, BlindGuard demonstrates competitive defense capability.
Preliminary
@@
In this paper, we focus on three types of primary attack modalities against MAS, i.e., prompt injection, memory poisoning, and tool exploitation (Wang et al. 2025). ❶ Prompt injection attacks manipulate agent outputs by corrupting either the system prompt $\mathcal{P}_{\text{sys}}$ or user inputs $\mathcal{P}_{\text{usr}}$, inducing malicious responses through carefully crafted textual perturbations. ❷ Memory poisoning targets the $\text{Mem}_{i}$ component by injecting fabricated interaction histories or poisoning external knowledge bases, thereby distorting the contextual understanding of the agent. ❸ Tool exploitation leverages vulnerabilities in external plugins ($\text{Plugin}_{i}$) to execute harmful operations such as unauthorized data access or privilege escalation. These attacks collectively transform the original system $\mathcal{G}$ into a compromised state $\widetilde{\mathcal{G}}$, where a subset of agents $\mathcal{V}_{\text{atk}}\subseteq\mathcal{V}$ exhibit adversarial behaviors while maintaining superficial operational normality.
@@
Supervised defense approaches leverage known attack patterns and labeled malicious samples to train detection models. Given a set of attacked MAS (with role and interaction description) where each MAS $\widetilde{\mathcal{G}}$ has labeled agents $\mathcal{V}=\mathcal{V}_{\text{norm}}\cup\mathcal{V}_{\text{mal}}$ where $\mathcal{V}_{\text{norm}}$ denotes normal agents and $\mathcal{V}_{\text{mal}}$ represents known malicious ones, the objective function typically minimizes:
$\mathcal{L}_{sup}=\sum_{v_{i}\in\mathcal{V}}\left[y_{i}\cdot f_{\theta}(\widetilde{\mathcal{G}},v_{i})+(1-y_{i})\cdot(1-f_{\theta}(\widetilde{\mathcal{G}},v_{i}))\right],$ (1)
where $y_{i}\in\{0,1\}$ indicates ground-truth labels (0 represents normal and 1 represents malicious) and $f_{\theta}:\mathbb{R}^{d}\rightarrow[0,1]$ is a classifier-based supervised GAD model parameterized by $\theta$. After training, the predicted anomaly scores of a given MAS are used to identify malicious agents $\mathcal{V}_{\text{atk}}$ for subsequent remediation, such as isolating malicious nodes or pruning suspicious communication links within the MAS.
A summary of related works is given in Appendix A.
@@
Methodology
While the defense approaches following the supervised paradigm show promising performance under controlled conditions, their reliance on labeled malicious agents hinders their applicability in real-world and MAS. To fill the gap, we propose a more practical unsupervised defense paradigm to extend the applicability of MAS safeguarding against any unknown attack. Based on the new paradigm, we proposed a novel approach, BlindGuard, which incorporates a specifically designed unsupervised GAD model that detects malicious agents without requiring any labeled data or prior knowledge of attack types as shown in Figure 2. In this section, we first formulate the unsupervised defense paradigm, and then introduce the core components of BlindGuard, i.e., hierarchical agent encoder, corruption-guided attack detector, and pruning-based remediation.
Unsupervised Defense Paradigm
@@
While other components in the unsupervised paradigm are similar to the supervised one, the central challenge is the architecture design and training strategy of the unsupervised detection model $f_{\theta}(\cdot,\cdot)$. Although existing unsupervised graph anomaly detection (GAD) methods (Pan et al. 2025, 2023; Ding et al. 2019; Qiao and Pang 2023; Liu et al. 2021; Li et al. 2024a) may serve as potential candidates, they are insufficient for the malicious agent identification task in MAS due to their limited capacity to capture multi-level agent interactions and their reliance on misaligned assumptions about anomaly patterns. Therefore, in BlindGuard, we introduce a specially designed unsupervised GAD model for malicious agent detection in MAS, with detailed descriptions provided in the following subsections.
Hierarchical Agent Encoder
@@
To process MAS with graph learning models, a key step is to convert the agent-level textual responses into node features. Given an agent $v_{i}$, the textual response $R_{i}$ is encoded into a pre-trained SentenceBERT (Reimers and Gurevych 2019) to map the response text to a dense vector $\mathbf{x}_{i}$:
@@
After acquiring the agent node features, we establish a GNN model in BlindGuard to learn expressive agent representations, which are subsequently used for malicious agent classification. While conventional GNNs (Wu et al. 2020; Kipf and Welling 2017) and GAD models (Qiao et al. 2024) are typically based on local neighborhood aggregation, they may overlook the global interaction patterns of the whole graph. Such global patterns, however, are essential for accurately detecting malicious agents in MAS, since malicious agents may coordinate their actions or influence others indirectly, requiring a system-level view to uncover these threats.
To bridge the gap, in BlindGuard, we design a hierarchical graph encoder that explicitly constructs the agent-level representations by incorporating information from three levels: ❶ Agent level, which captures individual semantic features of each agent derived from its textual response; ❷ Neighbor level, which aggregates contextual information from directly connected agents to model local interactions; and ❸ System level, which integrates global information across the entire MAS graph to capture long-range dependencies and collective behavior patterns. To implement this, we design a “summarization-transformation” architecture for multi-scale information fusion, similar to the “propagation-transformation” architecture of some lightweight GNNs (Wu et al. 2019; Zhang et al. 2022). Different from the propagation operations that only aggregate the 1-hop neighbors, in our summarization step, we integrate three complementary perspectives: ego-level features $\mathbf{h}_{i}^{\text{self}}$ to capture information of individual agent, neighbor-level features $\mathbf{h}_{i}^{\text{neigh}}$ to model local contexts, and global-level features $\mathbf{h}_{i}^{\text{graph}}$ to expose system-wide contexts. After the integration, we use a unified transformation to learn the compact representation for each agent. Formally, the representation $\mathbf{z}_{i}$ of agent $v_{i}$ can be calculated by:
$\mathbf{h}_{i}^{\text{self}}\!=\!\mathbf{x}_{i},\ \mathbf{h}_{i}^{\text{neigh}}\!=\!\sum_{j\in\mathcal{N}(i)}\hat{A}_{ij}\mathbf{x}_{j},\ \mathbf{h}_{i}^{\text{graph}}\!=\!\frac{1}{N}\sum_{k=1}^{N}\mathbf{x}_{k},$ (3)
$\mathbf{z}_{i}=g_{\theta}\left(\,\mathbf{h}_{i}^{\text{self}}\parallel\mathbf{h}_{i}^{\text{neigh}}\parallel\mathbf{h}_{i}^{\text{graph}}\,\right),$ (4)
@@
Corruption-Guided Attack Detector
Following the agent encoder, our corruption-guided attack detector aims to identify the malicious agents without any prior knowledge. Since ground-truth responses from attacked agents are unavailable during the training phase, we adopt a corruption-based strategy to simulate the semantic perturbations induced by adversarial attacks. Based on the simulated samples, we leverage a supervised contrastive learning objective to train the detection model, and then use a contextual similarity measurement to evaluate the abnormality of agents during inference.

### v2

@@ Abstract
The security of LLM-based multi-agent systems (MAS) is critically threatened by propagation vulnerability, where malicious agents can distort collective decision-making through inter-agent interactions. While existing supervised defense methods demonstrate promising performance, they may be impractical in real-world scenarios due to their heavy reliance on labeled malicious agents to train a supervised malicious detection model. To enable practical and generalizable MAS defenses, in this paper, we propose BlindGuard, an unsupervised defense method that learns without requiring any attack-specific labels or prior knowledge of malicious behaviors. To this end, we establish a hierarchical agent encoder to capture individual, neighborhood, and global interaction patterns of each agent, providing a comprehensive understanding for malicious agent detection. Meanwhile, we design a corruption-guided detector that consists of directional noise injection and contrastive learning, allowing effective detection model training solely on normal agent behaviors. Extensive experiments show that BlindGuard effectively detects diverse attack types across MAS with various communication patterns while maintaining superior generalizability compared to supervised baselines. The code is available at Code.
BlindGuard: Safeguarding LLM-based Multi-Agent Systems
under Unknown Attacks
Rui Miao1, Yixin Liu2, Yili Wang1, Xu Shen1, Yue Tan2, Yiwei Dai1, Shirui Pan2, Xin Wang1 1School of Artificial Intelligence, Jilin University, Changchun, China 2School of Information and Communication Technology, Griffith University, Gold Coast, Australia {miaorui24, shenxu23, daiyw23}@mails.jlu.edu.cn, {wangyili, xinwang}@jlu.edu.cn {yixin.liu, yue.tan, s.pan}@griffith.edu.au
footnotetext: Equal Contribution Corresponding Author
1 Introduction
@@
Rapid advancements in large language models (LLMs) have significantly improved their performance in various domains (Lei et al., 2024; Zheng et al., 2023; Yuan et al., 2024; Tian et al., 2025). By incorporating modular extensions such as memory (Tan et al., 2025), tool (Masterman et al., 2024), and role-playing (Kim et al., 2024), LLM-based autonomous agents have expanded their applicability, enabling more dynamic and interactive functionalities (Li et al., 2024b; Liu et al., 2026b). Building upon these advances, multi-agent systems (MAS) further amplify these benefits by facilitating collaborative interactions among specialized agents in more complex tasks (Guo et al., 2024). However, the increased reliance on inter-agent communication introduces additional risks in security, necessitating robust frameworks to safeguard sensitive data and regulate information flow.
@@
Figure 1: Comparison of supervised vs. unsupervised anomaly detection-based defense paradigms in MAS.
@@
Security studies (Andriushchenko et al., 2025; Gan et al., 2024; He et al., 2025) have identified significant vulnerabilities in external components of LLM-based agents. Beyond these risks at the single-agent level, misleading messages from a few malicious agents can propagate through collaborative reasoning, negatively impacting collective decisions (Yu et al., 2025). Such propagation vulnerability makes MAS susceptible to attacks such as prompt injection through compromised agents, misinformation propagation, and emergent malicious coordination Yu et al. (2024). To mitigate this, graph-based defense offers a promising solution by naturally modeling agent roles and interactions as a semantic structure. As demonstrated in Figure 1a, G-Safeguard Wang et al. (2025) integrates a detection-remediation framework, utilizing a supervised graph anomaly detection (GAD) model to identify malicious agents, followed by edge pruning to isolate their influence, thereby effectively safeguarding the MAS.
@@
However, the supervised GAD paradigm in defenses like G-Safeguard is often impractical, as it requires labeled malicious agents for training, which limits its availability and generalizability. First, adversarial attacks in real-world are often sparse and camouflaged, making it difficult to obtain well-annotated malicious agents for supervised training, limiting the availability of supervised GAD-based methods in deployments. Second, real-world MAS face diverse and evolving adversarial attacks, while binary GAD models are typically trained to detect a specific type of malicious behavior. Such the single-purpose design lacks the generalizability to detect novel or unseen patterns. These limitations raise a critical research question: Can we design a defense framework for MAS without relying on labeled attack agents?
@@
To answer the above question, as illustrated in Figure 1b, unsupervised GAD models (Ding et al., 2019; Ma et al., 2022; Li et al., 2024a; Liu et al., 2026a; Pan et al., 2025b) offer a promising solution forward by detecting anomalies using only normal MAS interaction data, potentially alleviating the limitations in availability and generalizability. However, methods not designed for MAS may lead to suboptimal performance due to the following gaps. Gap 1 - Limited multi-level contextual awareness: Identifying malicious agents requires integrating information across multiple levels, including individual behaviors, local neighborhoods, and global system. However, most existing GAD methods (Liu et al., 2021; Pan et al., 2023, 2025a, 2026b) primarily focus on local properties (e.g., local affinity), lacking the system-level understanding. Gap 2 - Misalignment of anomalous behavior assumptions: Most unsupervised GAD methods assume anomalies manifest through structural deviations (e.g., low homophily (Qiao and Pang, 2023) or rare connectivity patterns (Liu et al., 2021)). In contrast, malicious agents often exhibit semantic anomalies (such as deceptive intent (Yu et al., 2024)) that do not well match these assumptions.
@@
To fill the gaps, in this paper, we propose a novel defense method, termed BlindGuard, that can be trained without any labeled data or prior knowledge of attacks. BlindGuard employs an unsupervised GAD model to identify malicious agents, followed by edge pruning to suppress adversarial propagation. To bridge Gap 1, we introduce a hierarchical encoder to incorporate the information of individual agent features, local neighborhood aggregation, and global system context simultaneously. As a result, the encoder captures comprehensive representations of agents to support malicious agent detection. To mitigate Gap 2, we propose a corruption-guided attack detector for agent abnormality estimation. We simulate the malicious behaviors via semantic-level corruption, which is utilized to optimize the detector via a supervised contrastive learning objective. The training of BlindGuard only requires a small amount of normal MAS interaction data, and the learned model can generalize to various types of attacks. To sum up, the contributions of this paper are three-fold:
@@
  • Scenario. We investigate the scenario of MAS safeguarding without relying on labeled data or prior knowledge of attacks, which is more practical and applicable to real-world MAS.
@@
  • Experiments. We extensively evaluate BlindGuard under rigorous real-world conditions. Through comprehensive testing on 4 MAS structures with 3 attack strategies, BlindGuard demonstrates competitive performance.
2 Preliminary
@@
In this paper, we focus on three types of primary attack modalities against MAS, i.e., prompt injection, memory poisoning, and tool exploitation (Wang et al., 2025). ❶ Prompt injection attacks manipulate agent outputs by corrupting either the system prompt $\mathcal{P}_{\text{sys}}$ or user inputs $\mathcal{P}_{\text{usr}}$, inducing malicious responses through carefully crafted textual perturbations. ❷ Memory poisoning targets the $\text{Mem}_{i}$ component by injecting fabricated interaction histories or poisoning external knowledge bases, thereby distorting the contextual understanding of the agent. ❸ Tool exploitation leverages vulnerabilities in external plugins ($\text{Plugin}_{i}$) to execute harmful operations such as unauthorized data access or privilege escalation. These attacks transform the original system $\mathcal{G}$ into a compromised state $\widetilde{\mathcal{G}}$, where a subset of agents $\mathcal{V}_{\text{atk}}\subseteq\mathcal{V}$ exhibit adversarial behaviors while maintaining superficial operational normality.
@@
Supervised defense approaches leverage known attack patterns and labeled malicious samples to train detection models. Given a set of attacked MAS (with role and interaction description) where each MAS $\widetilde{\mathcal{G}}$ has labeled agents $\mathcal{V}=\mathcal{V}_{\text{norm}}\cup\mathcal{V}_{\text{mal}}$ where $\mathcal{V}_{\text{norm}}$ denotes normal agents and $\mathcal{V}_{\text{mal}}$ represents known malicious ones, the objective typically minimizes:
$\mathcal{L}_{sup}=\sum_{v_{i}\in\mathcal{V}}\ell_{CE}\big(y_{i},f_{\theta}(\widetilde{\mathcal{G}},v_{i})\big),$ (1)
where $y_{i}\in\{0,1\}$ indicates ground-truth labels (0 represents normal and 1 represents malicious) and $f_{\theta}:\mathbb{R}^{d}\rightarrow[0,1]$ is a classifier-based supervised GAD model parameterized by $\theta$. After training, the predicted anomaly scores of a given MAS are used to identify malicious agents $\mathcal{V}_{\text{atk}}$ for subsequent remediation, such as isolating malicious nodes or pruning suspicious communication links. A summary of related works is given in Appendix A.
@@
3 Methodology
While the defense approaches following the supervised paradigm show promising performance under controlled conditions, their reliance on labeled malicious agents hinders their applicability in real-world and MAS. To fill the gap, we propose a more practical unsupervised defense paradigm to extend the applicability of MAS safeguarding against any unknown attack. Based on the new paradigm, we proposed a novel approach, BlindGuard, which incorporates a specifically designed unsupervised GAD model that detects malicious agents without requiring any labeled data or prior knowledge of attacks as shown in Figure 2. In this section, we first formulate the unsupervised defense paradigm, and then introduce the core components of BlindGuard, i.e., hierarchical agent encoder, corruption-guided attack detector, and pruning-based remediation.
3.1 Unsupervised Defense Paradigm
@@
While other components in the unsupervised paradigm are similar to the supervised one, the central challenge is the architecture design and training strategy of the unsupervised detection model $f_{\theta}(\cdot,\cdot)$. Although existing unsupervised graph anomaly detection (GAD) methods may serve as potential candidates, they are insufficient for the malicious agent identification task in MAS due to their limited capacity to capture multi-level agent interactions and their reliance on misaligned assumptions about anomaly patterns. Therefore, in BlindGuard, we introduce a specially designed unsupervised GAD model for malicious agent detection in MAS, with detailed descriptions provided in the following subsections.
3.2 Hierarchical Agent Encoder
@@
To process MAS with graph learning models, a key step is to convert the agent-level textual responses into node features. Given an agent $v_{i}$, the textual response $R_{i}$ is encoded into a pre-trained SentenceBERT Reimers and Gurevych (2019) to map the response text to a dense vector $\mathbf{x}_{i}$:
@@
After acquiring the agent node features, we establish a GNN model in BlindGuard to learn expressive agent representations, which are subsequently used for malicious agent classification. While conventional GNNs Kipf and Welling (2017); Wu et al. (2020); Miao et al. (2024) and GAD models Qiao et al. (2024) are typically based on local neighborhood aggregation, they may overlook the global interaction patterns of the whole graph. Such global patterns, however, are essential for accurately detecting malicious agents in MAS, since malicious agents may coordinate their actions or influence others indirectly, requiring a system-level view to uncover these threats.
To bridge the gap, in BlindGuard, we design a hierarchical graph encoder that explicitly constructs the agent-level representations by incorporating information from three levels: ❶ Agent level, which captures individual semantic features of each agent derived from its textual response; ❷ Neighbor level, which aggregates contextual information from directly connected agents to model local interactions; and ❸ System level, which integrates global information across the entire MAS graph to capture long-range dependencies and collective behavior patterns. To implement this, we design a “summarization-transformation” architecture for multi-scale information fusion, similar to the “propagation-transformation” architecture of some lightweight GNNs Wu et al. (2019); Zhang et al. (2022). Different from the propagation operations that only aggregate the 1-hop neighbors, in our summarization step, we integrate three complementary perspectives: ego-level features $\mathbf{h}_{i}^{\text{s}}$ to capture information of individual agent, neighbor-level features $\mathbf{h}_{i}^{\text{n}}$ to model local contexts, and global-level features $\mathbf{h}_{i}^{\text{g}}$ to expose system-wide contexts. After the integration, we use a unified transformation to learn the compact representation for each agent. Formally, the representation $\mathbf{z}_{i}$ of agent $v_{i}$ can be calculated by:
$\mathbf{h}_{i}^{\text{s}}\!=\!\mathbf{x}_{i},\ \mathbf{h}_{i}^{\text{n}}\!=\!\sum_{j\in\mathcal{N}(i)}\hat{A}_{ij}\mathbf{x}_{j},\ \mathbf{h}_{i}^{\text{g}}\!=\!\frac{1}{N}\sum_{k=1}^{N}\mathbf{x}_{k},$ (3)
$\mathbf{z}_{i}=g_{\theta}\left(\,\mathbf{h}_{i}^{\text{s}}\parallel\mathbf{h}_{i}^{\text{n}}\parallel\mathbf{h}_{i}^{\text{g}}\,\right),$ (4)
@@
3.3 Corruption-Guided Attack Detector
Following the agent encoder, our corruption-guided attack detector aims to identify the malicious agents without any prior knowledge. Since ground-truth responses from attacked agents are unavailable during the training phase, we adopt a corruption-based strategy to simulate the semantic perturbations induced by adversarial attacks. Based on the simulated samples, we leverage a supervised contrastive learning objective to train the detection model, and use a contextual similarity measurement to evaluate the abnormality of agents during inference.